OpenStudy (anonymous):

Expressing a.) cos(3θ), b.) sin(4θ) in terms of cos(θ) and sin(θ).

OpenStudy (campbell_st):

well as a start for the (a) use the sum of 2 angles \[\cos(3\theta) = \cos(2\theta + \theta)\] for (b) use \[\sin(4 \theta) = \sin(2 \theta + 2 \theta) \] for both questions it should be reasonably straight forward from here

OpenStudy (anonymous):

How it goes when i use De Moivre's formula?

OpenStudy (campbell_st):

I wouldn't have thought you needed De Moivre...and thats used in complex numbers

OpenStudy (anonymous):

sry, i forgot to mention those in my question.